# Neural Video Fields Editing

### 12 Dec 2023 v1

@@ 4.3 Application
Here we demonstrate some practical applications of NVEdit to establish that it not only enables multiple editing types (e.g., changing shape, editing scene, and style transfer, etc.), but also accomplishes frame interpolation without any fine-tuning or post-processing operations.
Consistent video editing. Fig. 1 has displayed our results on various scenes and instructions, illustrating the temporal consistency and adherence to instructions achieved by our edits. On the left, we replace the wolf with other animals. The added animals maintain the same posture and motion as the wolf. Crucially, the background remains unchanged, whilst inter-frame variation is smooth and consistent throughout the edited video. On the right, we showcase the effects of scene and style editing. Our editing strictly follows the instructions, generating temporal consistent videos while accurately inheriting original features. In addition, due to the impressive encoding efficiency, NVEdit enables long video editing. We present our results on long videos comprising 600 frames in Fig. 5, whose resolution is 960 $\times$ 544. Again, different types of editing are applied to these videos. One can see that our method faithfully edits frames according to the user-given instructions, preserving information that does not require modification, such as motion and semantic layout, etc. Inter-frame consistency is also well-maintained after the field editing stage, allowing NVEdit to render editing effects without unnatural flickers.
[…]
Figure 8: Results of NVEdit trained w/ and w/o our Auxiliary Mask (AM). Without AM, the background is incorrectly changed, while our final version achieves precise local editing.
@@ 4.4 Ablation Study
We mainly conduct an ablation study on the proposed Auxiliary Mask (AM), and substitute IP2P+ with an alternative image processing algorithm. The former demonstrates our improvement potential, while the latter validates the replacement ability for the T2I model within our framework.
Ablation on AM. As we have showcased AM and the results of IP2P and IP2P+ in Fig. 3, here we provide the rendered frames of the same instruction from NVEdit trained w/ and w/o the AM to underscore its importance. Since AM points out the editable regions for local editing, IP2P+ enables more precise editing and produces higher-quality pseudo-GTs to optimize our NVF. Therefore, as depicted in Fig. 8, NVEdit outperforms the ablation setting, which exhibits a clearer face and an unchanged background.
Substitute the T2I model. Beyond IP2P, various image processing algorithms can be applied in NVEdit for diverse requirements. Here we take R-ESRGAN [44] as an example. By utilizing it to conduct image super-resolution on each rendered frame and produce pseudo-GT, our method seamlessly enhances the resolution of videos. Illustrated in Fig. 9, NVEdit initially fits a low-resolution video (top row), then optimized with pseudo-high-resolution ($\times$4) GTs, achieving video super-resolution (bottom row).
Figure 9: Video super-resolution results ($\times$4) achieved by using an image super-resolution algorithm [44] to optimize the video field.
@@ 5 Conclusion
This paper presents a memory-efficient framework NVEdit to edit videos with off-the-shelf image processing methods. We design a neural video field to enable long video editing and learn temporal priors for inter-frame consistency. In order to impart editing effects to the trained video field while preserving its temporal priors, a progressive optimization strategy is applied. Furthermore, to improve editing quality, we upgrade IP2P with the proposed auxiliary mask to benefit its precision in local editing, showcasing remarkable control over editable regions without additional computation, dubbed IP2P+. NVEdit outperforms existing baselines, demonstrating substantial improvements in local editing, temporal consistency, and the length of editable video. At last, we shed the unique advantages and immense potential of neural representations in video editing, illustrating how they can be synergized with T2I models. Both the video field and IP2P+ can be replaced or improved for various requirements, which inspires future research.
@@
Limitations. Although we have achieved SOTA consistency in editing, temporal priors are still inevitably affected to some extent in the field editing stage. Moreover, the increased frames require more iterations in the field editing stage, resulting in more time costs. We suppose that, in the future, only editing keyframes and using frame interpolation of NVF to propagate the editing effect throughout the entire video is a potential acceleration solution.
[…]
Supplementary Material
Algorithm 1 NVEdit Algorithm
Require:
- $\widetilde{\boldsymbol{\theta}}$: Parameters of the desired video field $f(\widetilde{\boldsymbol{\theta}})$
- $Edit$: The chosen T2I model
Input:
- V: Source video
- $P$: User-given instruction
$\triangleright$ The video fitting stage
while not converge do
       $\widetilde{\textbf{V}}[x,y,t]=f_{\widetilde{\boldsymbol{\theta}}}(x,y,t)$
       $\widetilde{\boldsymbol{\theta}}\leftarrow\operatorname{argmin}_{\widetilde{\boldsymbol{\theta}}}\!\!\!\!\!\underset{(x,y,t)\in\boldsymbol{\psi}}{\sum}\!\!\!\!\!\!\!\|\textbf{V}[x,y,t]-\widetilde{\textbf{V}}[x,y,t]\|_{2}^{2}$$\triangleright$ $\mathcal{L}_{\rm{fit}}(\widetilde{\boldsymbol{\theta}})$
end while
$\triangleright$ The field editing stage
while not converge do
       $\widetilde{\textbf{F}}_{t}=f_{\widetilde{\boldsymbol{\theta}}}(\textbf{\emph{X}},\textbf{\emph{Y}},t)$
       $\overline{\textbf{F}}_{t}=Edit(\widetilde{\textbf{F}}_{t},\textbf{V}[\textbf{\emph{X}},\textbf{\emph{Y}},t],P)$triangleright Edit the frame
       $\widetilde{\boldsymbol{\theta}}\leftarrow\operatorname{argmin}_{\widetilde{\boldsymbol{\theta}}}\|\overline{\textbf{F}}_{t}-\widetilde{\textbf{F}}_{t}\|_{1}$triangleright $\mathcal{L}_{\rm{opt}}(\widetilde{\boldsymbol{\theta}})$
end while
$\triangleright$ Inference
[…]
       $\widetilde{\textbf{V}}=f_{\widetilde{\boldsymbol{\theta}}}(\textbf{\emph{X}},\textbf{\emph{Y}},\textbf{\emph{T}})$triangleright Eq. 6 in main paper
       return $\widetilde{\textbf{V}}$
end function
Output:
- $\widetilde{\textbf{V}}$: Final edited video
6 Implementation Details
Pseudo algorithm code. Our full algorithm is shown in Algorithm 1, which includes the process of video fitting, field editing, and inference. NVEdit first learns motion and temporal priors of the given video (video fitting stage). Then a T2I model is applied to edit the rendered frame $\widetilde{\textbf{F}}_{t}$, which is utilized to update the neural video field for editing effects (field editing stage). Here we adopt Instruct-Pix2Pix+ as the function $Edit$. Finally, the edited video can be rendered through the trained parameters.
Hyperparameters. We mainly introduce three hyperparameters in the proposed method:
- ${t}_{u}\in[1,T]$: The upper boundary for the initial sampling timestep in Instruct-Pix2Pix+. A smaller value results in a diminished editing amplitude, whereas a larger value corresponds to a more pronounced editing effect.
- ${t}_{l}\in[1,T]$: The lower limit for the initial sampling timestep in Instruct-Pix2Pix+. Bigger ${t}_{l}$ correlates with a greater edit.
- (Optional) $\tau\in[0,1]$: The threshold for our auxiliary mask used in Instruct-Pix2Pix+, which is only adopted in the case of local editing. Bigger $\tau$ enables larger editable region.
In the context of global editing, we set ${t}_{u}=0.98T$ and ${t}_{l}=0.42T$, with the omission of an auxiliary mask. For local editing, we set ${t}_{u}=0.98T$ and ${t}_{l}=0.82T$. $\tau=0.1$ is an empirical value, which has been able to perform well in most applications and users can adjust it slightly based on their specific requirements.
Figure 10: Visual results of the composed editing. We recurrently run the field editing stage with different instructions to achieve blended editing effects.
Figure 11: Four sets of visual results from NVEdit, which enables multiple types of editing, including style transfer and shape variation. The instructions for editing are noted under each set.
Figure 12: Editing results of a long video. The frame number is provided below each image (from $f_{1}$ to $f_{50}$.)
7 More Visual Results
You can further view more editing results in our project. In this section, we succinctly introduce an extension work (composed editing) and furnish additional subjective results to substantiate the efficacy of NVEdit. All results provided in this section are available in our project.
Composed editing. In the main paper, we have implemented text-driven editing based on the user-given instructions. Furthermore, we attempt to apply field editing stage recurrently with different instructions to achieve composed editing. As depicted in Fig. 10, the image in the third column simultaneously satisfies both instructions, validating the feasibility of composed editing achieved through the recurrent invocation of the field editing stage.
More editing results. NVEdit enables multiple types of editing, including style transfer and shape variation. To further validate its superior editing capabilities, in Fig. 11, we provide additional four sets of editing results. Each set comprises the original video displayed in the upper line, juxtaposed with the corresponding edited results in the lower line. We provide their related editing instructions under the edited videos of each set.
Long video editing. Notice that NVEdit enables high-quality editing for long videos, exhibiting a significant practical merit. For a more intuitive validation, detailed results are elucidated in Fig. 12, which includes a 50-frame original video juxtaposed with its edited version. We turn the running horse into a donkey and the corresponding frame numbers (from $f_{1}$ to $f_{50}$) are annotated beneath their respective outcomes. More long video editing results can be accessed in our project.
Method Runtime (s)
4 8 12 16
TokenFlow 126.51 211.8 283.24 404.19
Video-P2P 212.82 367.06 512.38 677.37
V2V-zero 116.05 235.05 371.26 518.29
Fatezero 76.67 129.41 175.19 224.77
CoDeF 256.34 397.61 649.62 997.99
Ours 428.79 609.72 905.97 1104.26
Table 2: The comparison of different methods in terms of runtime costs. Experiments are conducted on a single NVIDIA A40 GPU, using videos at a resolution of 480 $\times$ 480.
8 Comparison on Runtime
Due to the optimization strategy in the field editing stage, our training time scales linearly with the duration of the videos to be edited, as discussed in the limitation part of the main paper. We provide a detailed quantitative comparison in Tab. 2, which compares the runtime of different methods on videos with different frame numbers, including 4, 8, 12, and 16. Their resolution is all 480 $\times$ 480. One can see that our time cost is slightly higher than that of the compared SOTA methods. It is important to note, however, that these SOTA methods face constraints in editing lengthy videos due to finite GPU memory, whereas NVEdit successfully overcomes this limitation, as illustrated in Fig. 12. To accelerate editing, we suppose that instead of sampling frames across the entire video, asign key frames in advance and only edit them for supervision in the field editing stage is a feasible solution.
9 Limitation and Future Work
We impart editing effects through the field editing stage, which makes our performance to be constrained by the effectiveness of the chosen T2I model. To enhance the editing effects, we applied IP2P+ in the main paper. Notice that NVEdit also supports other image editing or processing algorithms, which enables different editing effects and even various downstream video tasks, e.g., the adoption of R-ESRGAN in Sec. 4.4 of the main paper for video super-resolution. We leave further research to future work.
Figure 13: Some visual samples of the original image, the results of IP2P, the proposed auxiliary mask, and results of IP2P+.
10 Results of Instruct-Pix2Pix+
In this paper, we develop a competitive text-driven image editing algorithm, called Instruct-Pix2Pix+ (IP2P+), which achieves impressive effects in local editing without any additional computation. Details have been discussed in Sec. 3.3 of the main paper. Here we additionally provide some results of this valuable image editing approach in Fig. 13. Hoping that this innovation inspires future artificial intelligence generative content work.

### 09 Mar 2024 v2

@@ 4.3 Application
Here we demonstrate some other practical applications of NVEdit to establish that it not only enables multiple editing types (e.g., changing shape, editing scene, and style transfer, etc.), but also accomplishes frame interpolation without any fine-tuning or post-processing operations.
Consistent video editing. Fig. 1 has displayed our results on various scenes and instructions, illustrating the temporal consistency and adherence to instructions achieved by our edits. On the left, we replace the wolf with other animals. The added animals maintain the same posture and motion as the wolf. Crucially, the background remains unchanged, whilst inter-frame variation is smooth and consistent throughout the edited video. On the right, we showcase the effects of scene and style editing. Our editing strictly follows the instructions, generating temporal consistent videos while accurately inheriting original features. In addition, due to the impressive encoding efficiency, NVEdit enables long video editing. We present our results on long videos comprising 600 frames in Fig. 6, whose resolution is 960 $\times$ 544. Again, different types of editing are applied to these videos. One can see that our method faithfully edits frames according to the user-given instructions, while preserving information that does not require modification, such as motion and semantic layout, etc. Inter-frame consistency is also well-maintained after the field editing stage, allowing NVEdit to render editing effects without unnatural flickers. We present video results in the supplementary material for a more intuitive illustration.
Table 2: The quantitative study on other applications of NVEdit.
Table 3: Ablation study on progressive optimization. IP2P+ serves as a control group.
Metrics PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Interpolation 33.81 0.961 0.302
SR ($\times$4) 28.95 0.812 0.246
Metrics Tem-Con $\uparrow$ Fram-Acc $\uparrow$ Vid-Score $\uparrow$
IP2P+ 0.966 0.957 0.309
Fix $s_{P}$ 0.959 0.960 0.313
Table 3: Ablation study on progressive optimization. IP2P+ serves as a control group.
Frame interpolation. Since the neural video field encodes a video as a temporally continuous signal [8, 22], we can seamlessly extend the learned signals along the temporal axis, i.e., frame interpolation, without any additional operations. We validate its ability of frame interpolation in Tab. 3 with the UVG-HD dataset [27]. Three well-known objective metrics are used for evaluation, including PSNR, SSIM [51], and LPIPS [57]. We evenly select 300 frames from 600 frames every other frame and calculate average metrics of our interpolated 300 frames with the Ground Truth. Results in the first row of Tab. 3 prove that NVF successfully predicts intermediate frames with competitive accuracy.
Figure 7: Results of NVEdit trained with (w/) and without (w/o) our proposed Auxiliary Mask (AM). Without AM, the background is incorrectly changed, while our final version achieves precise local editing.
@@ 4.4 Ablation Study
In this section, we conduct ablation research on progressive optimization in the field editing stage, Auxiliary Mask (AM) used in IP2P+, and the adopted T2I model. The first study aims to prove the importance of progressive training in generating temporal consistent videos, and the second demonstrates the potential of our method to improve video editing performance by utilizing new image related technologies. Finally, we replace IP2P+ with another image processing algorithm to validate the substitutability of the T2I model within our framework.
Ablation on progressive optimization. To investigate the contribution of progressive training, we first fix the random seed of IP2P+ and apply it per-frame, with the results as the control group. Besides, we train an NVEdit but fix the value of $s_{P}$ in the field editing stage. Results are reported in Tab. 3, in which we adopt the same objective metrics used in Tab. 1. One can see that compared to directly using IP2P+, training NVEdit with a fixed $s_{P}$ actually leads to significant degradation on ‘Tem-Con’, a measure of inter frame consistency. We believe this is because the fixed editing intensity disrupts the prior learned in the video fitting stage. In contrast, as listed in Tab. 1, the complete NVEdit achieves optimal results across all three metrics compared to these two ablation settings, which proves the effectiveness of our progressive optimization strategy.
Ablation on AM. We have showcased AM and the results of IP2P and IP2P+ in Fig. 3, here we provide the rendered frames of the same instruction from NVEdit trained w/ and w/o the AM to underscore its importance. Since AM points out the editable regions for local editing, IP2P+ enables more precise editing and produces higher-quality pseudo-GTs to optimize our NVF. Therefore, as depicted in Fig. 7, NVEdit outperforms the ablation setting, which exhibits a clearer face and an unchanged background. We further provide more visual comparisons between IP2P and IP2P+ in the supplementary material.
Ablation on the T2I model. Beyond IP2P, various image processing algorithms can be also applied in NVEdit for diverse requirements. Here we take R-ESRGAN [48] as an example. By utilizing it to conduct image super-resolution on each rendered frame and produce pseudo-GT, our method seamlessly enhances the resolution of videos. As illustrated in Fig. 8, we select two original frames and provide their recovered versions on the right. To achieve it, in the video fitting stage, NVEdit initially fits the original low-resolution video, then is optimized with the pseudo-high-resolution ($\times$4) GTs in the field editing stage. The final rendered frames are listed on the right for comparison. Besides, we further calculate quantitative results on UVG-HD dataset and report them in the second row of Tab. 3. Videos are first downsampled by four times and then recovered. These superior results validate the potential of NVEdit that adopting image algorithms to achieve different video tasks in a plug-and-play manner.
Figure 8: Video super-resolution results ($\times$4) generated by using an image super-resolution algorithm [48] to optimize the video field. Two original low-resolution frames and their recovered versions are provided for intuitive illustration.
@@ 5 Conclusion
This paper presents a memory-efficient framework NVEdit to edit videos with off-the-shelf image processing methods. We design a neural video field to enable long video editing and learn temporal priors for inter-frame consistency. In order to impart editing effects to the trained video field while preserving its temporal priors, a progressive optimization strategy is developed. Furthermore, to improve editing quality, we upgrade IP2P with the proposed auxiliary mask to benefit its precision in local editing, showcasing remarkable control over editable regions without additional computation, dubbed IP2P+. NVEdit outperforms existing baselines, demonstrating substantial improvements in local editing, temporal consistency, and editable video length. At last, we shed the unique advantages and immense potential of neural representations in video editing, illustrating how they can be synergized with T2I models. Both the video field and IP2P+ can be replaced or improved for various requirements, which inspires future research.
@@
Limitations. Although we have achieved SOTA consistency in editing, temporal priors are still inevitably affected to some extent in the field editing stage. Moreover, the increased frames require more iterations in the field editing stage, resulting in more time costs. We believe that, in the future, only editing key frames and using frame interpolation of NVF to propagate the editing effect throughout the entire video is a potential acceleration solution.